A sample of oxygen occupies a volume of 45ml at 91 Degrees Celsius. What will be the volume of oxygen when temperature drops at 10 Degrees Celsius


1
Expert's answer
2021-09-27T06:33:24-0400

V1/T1 = V2/T2

45/364 = V2/283

V2= 34.99ml